OpenStudy (anonymous):

simplify the expression (-8i)(2i)

OpenStudy (anonymous):

please?

OpenStudy (ash2326):

\[-8i \times 2i\] \[-16 \times i \times i\] \[-16 \times i^2\] We know i^2=-1 as \(i=\sqrt{-1}\) so we get \[-16 \times -1\] I think you can simplify the rest